OpenStudy (anonymous):

What is the y coordinate of the y−intercept of the line that passes through the points (−4, −1) and (4, 5)? Numerical Answers Expected! need helppp please...

OpenStudy (helder_edwin):

first find the slope do u know how?

OpenStudy (anonymous):

yes and then

OpenStudy (helder_edwin):

do it! tell me what u get

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i get 6/8

OpenStudy (helder_edwin):

yes reducing u get 3/4 right?

OpenStudy (anonymous):

oh yeah

OpenStudy (helder_edwin):

now u have one point of the line (4,5) or (-4,-1) whichever (but not both) and u have m=3/4 the slope the line has the following equation \[ \large y-y_0=m(x-x_0) \] where (x0,y0) is one of the points u have plug in and tell me what u get

OpenStudy (anonymous):

y-5 = 3/4(x-4) im not sure what i have to do :/

OpenStudy (helder_edwin):

great this is the equation of the line now the y-intercept is what u get when u have x=0

OpenStudy (helder_edwin):

so \[ \large y-5=\frac{3}{4}(0-4) \] \[ \large y=5+\frac{3}{4}(-4)= \]

OpenStudy (helder_edwin):

u finish it

OpenStudy (helder_edwin):

do u understand what i did?

OpenStudy (anonymous):

so the answer is 5

OpenStudy (helder_edwin):

no

OpenStudy (helder_edwin):

simplify the 4's what is left?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

no 2

OpenStudy (helder_edwin):

no \[ \large \frac{-4}{4}=-1 \] right=

OpenStudy (helder_edwin):

yes y=2

OpenStudy (helder_edwin):

heres is another way \[ \large y-5=\frac{3}{4}(x-4) \] \[ \large y-5=\frac{3}{4}x-3 \] \[ \large y=\frac{3}{4}x-3+5 \] \[ \large y=\frac{3}{4}x+2 \] this equation has the form \[ \large y=mx+b \] where m is the slope as before and b is the y-intercept

OpenStudy (anonymous):

ok thanks again.

OpenStudy (helder_edwin):

u r welcome